The product of two consecutive positive even integers is 14 more than their sum. Find the integers.
Answer by checkley77(12844)

X(X+2)=X+X+2+14
X^2+2X=2X+16
X^2-16=0
(X+4)(X-4)=0
X+4=0
X=-4 ans.
X-4=0
X=4 ans.
PROOF:
4*6=4+6+14
24=10+14
24=24